# Towards Galactic Archaeology with Inferred Ages of Giant Stars From Gaia Spectra

### v1

@@ 3.2 Age inference from RVS spectra using SIDRA-RVS
Fig. 4 is a SHapley Additive exPlanations (SHAP, Lundberg & Lee, 2017) bee-swarm plot of our model’s results shown in Fig. 2 and demonstrates the relative importance of the flux values at the different wavelength input features in the RVS spectra and how they impact the stellar-age model output. Fig. 4 shows the distribution of the impact across the top 10 high-impact wavelength pixel values. This visualisation is important for understanding our SIDRA-RVS model’s behaviour and the significance of different elemental lines. The x-axis represents the SHAP values, which indicate the impact of each wavelength input feature (normalised flux) on the prediction. SHAP values to the right of the centerline (positive values) suggest that the corresponding input feature contributes to a higher predicted age, whereas values to the left (negative values) indicate that the input feature results in a lower predicted age. The colour of the dots represents the relative values of the input features, with red indicating higher values and blue indicating lower values. Thus, if blue dots appear on the right side, it suggests that lower input feature values correspond to higher output values. The y-axis represents the features sorted by their importance from top to bottom. Here, the importance of the input feature is measured as the summation of the absolute values of SHAP value for all the data. The top elemental features, TiI: 869.47 nm, TiI: 869.46 nm (2nd), and TiI: 869.48 nm (6th), belong to the same absorption line. Similarly, TiI: 852.06 nm (3rd), TiI: 852.05 nm (8th), and TiI: 852.04 nm (10th) are part of the same line. Additionally, NI: 868.53 nm (4th) and NI: 858.54 nm (5th) are also in the same absorption line, making it the third most important line. Blue dots on the high SHAP value side indicate that a lower flux value at the TiI lines, representing a stronger TiI absorption feature in the spectra, predicts a higher age. On the other hand, red dots on the low SHAP value side suggest that a higher flux value at TiI, representing a weaker TiI absorption feature in the spectra, predicts a lower age.
Figure 5: Example of a high signal-to-noise ratio (SNR) spectrum of an observed star from Gaia DR3 (DR3 Star ID-1222988540219279360) shown in two wavelength regions: 846-858 nm (top panel) and 858-870 nm (bottom panel). The red dots on the spectrum represent the top 10 highest SHAP values from our SIDRA-RVS model results, overlaid with atomic lines from Recio-Blanco et al. (2023). The width of the atomic lines corresponds to its uncertainty. The vertical blue lines at 865.09 in the bottom panel indicate SiI line identified by Contursi et al. (2021). The stellar parameters, $\log_{10}{(T_{\mathrm{eff}})}$, $\log~{}g$, $[\mathrm{\alpha/M}]$ are obtained from APOGEE.
@@ 4.1 SIDRA-XP’s methodology
The relationships between stellar ages and stellar parameters, such as effective temperature, $T_{\mathrm{eff}}$, surface gravity, log $g$, and various abundance ratios, such as [C/Fe], [Fe/H], [N/Fe] and [$\mathrm{\alpha/M}$], can be complex and non-linear. ANNs are well-suited for this task because they can autonomously identify and model complex relationships without the need for explicitly programmed rules. ANNs are computational models inspired by the human brain’s neural networks. They consist of layers of interconnected nodes, or "neurons", which process input data to learn patterns and relationships. In supervised learning, the network’s parameters, such as the weights connecting the neurons, are trained and optimised to accurately reproduce known input-output pairs from a training set. Once trained, the ANN can efficiently predict the ages of the other stars based on their observed properties, using significantly less computational effort than the training process.
We split the data 80% training data and 20% testing data. To mitigate the data imbalance between young and old stars, we apply the same random-sampling data-augmentation technique described in Sec. 3.1 to address the underrepresented ages for the training set of the BINGO-FS24. For the data augmentation, we sample $T_{\mathrm{eff}}$, log $g$, [C/Fe], [Fe/H], [N/Fe], [$\alpha$/Fe] and age from the original star data using the random Gaussian based on the uncertainties in FS24 and BINGO.
@@ 4.2 Age inference from XP stellar parameters using SIDRA-XP
Figure 6: SIDRA-XP age predictions, $\mathrm{log}_{10}(\tau_{\mathrm{SIDRA}}$ $\mathrm{[Gyr]})$, versus the target BINGO age estimation in $\mathrm{log}_{10}(\tau_{\mathrm{BINGO}}$ $\mathrm{[Gyr]})$ for the XP parameters testing set colour-coded by [Fe/H] (left panel) and [$\mathrm{\alpha/M}$] (right panel) obtained from FS24. The upper panels show the predictions versus the target and the black line indicates the identity line. The lower panels represent the residuals between the SIDRA-XP’s $\mathrm{log}_{10}(\tau_{\mathrm{SIDRA}}$ $\mathrm{[Gyr]})$ prediction and BINGO’s true $\mathrm{log}_{10}(\tau_{\mathrm{BINGO}}$ $\mathrm{[Gyr]})$ denoted as $\mathrm{log}_{10}(\frac{\tau_{\mathrm{SIDRA}}}{\tau_{\mathrm{BINGO}}})$. The black filled circles and vertical error bars indicate the mean and the standard deviation of the residuals at the different $\mathrm{log}_{10}(\tau_{\mathrm{BINGO}}$ $\mathrm{[Gyr]})$ bins, respectively.
@@
The colours of the dots in Fig. 6 indicate [Fe/H] (left panel) and [$\mathrm{\alpha/M}$] (right panel). The figure shows that there is scatter in [Fe/H] and [$\alpha$/M] for the stars with similar age. This means that the SIDRA-XP model does not exclusively capture the correlations between the [$\mathrm{\alpha/M}$] or [Fe/H] and stellar age when using XP parameters, ensuring that the model’s learning process goes beyond these specific trends, capturing a broader range of stellar characteristics.
@@
Fig. 7 is a SHAP bee-swarm plot that illustrates the contribution of various input features to the predicted ages of stars using SIDRA-XP. Each dot represents an individual star in the test set result as shown in Fig. 6. The x-axis indicates the SHAP value, representing the impact of each feature on the model’s age prediction for a given star. The features are listed on the y-axis in order from the highest to lowest importance of the input features, which includes effective temperature, $T_{\mathrm{eff}}$, surface gravity, log $g$, and different abundance ratios, such as [C/Fe], [Fe/H], [N/Fe] and [$\mathrm{\alpha/M}$]. The distribution and spread of the dots indicate how strongly each XP stellar parameter feature influences the stellar age prediction, with dots further from zero having a more significant impact. The alignment of the dots along the x-axis shows whether the features have a positive or negative contribution to the predicted age. In the case of $T_{\mathrm{eff}}$, the bluer dots in the high SHAP values for $T_{\mathrm{eff}}$ indicate that low $T_{\mathrm{eff}}$ stars push the age prediction higher and the redder dots in the lower SHAP values mean that high $T_{\mathrm{eff}}$ stars push the age prediction to lower. This trend is consistent with with the right panel of Fig. 1 and what is found for SIDRA-RVS in Section 3.2.
@@
The [Fe/H] trend of Fig. 7 is puzzling. It shows the higher [Fe/H] leading to older age. However, there are some stars with blue (low [Fe/H]) in the positive SHAP value (older age). Hence, there seem to be a mixed trend of age-[Fe/H] relation. This likely reflects the flat trend of age-[Fe/H] relation in the thin disc, which is also discussed in the next section. On the other hand, the [$\alpha$/M] shows a clear trend of higher [$\alpha$/M] leading to older age. This is consistent with the known age-[$\alpha$/M] relation.
Our results demonstrate that the stellar parameters and [C/N] abundances derived from the XP spectra by FS24 provide reliable stellar age information for the giant stars. Unlike SIDRA-RVS, SIDRA-XP learned age-[C/N] relation in addition to the age dependence on $T_{\mathrm{eff}}$, $\log~{}g$ and [$\alpha$/M]. Also, SIDRA-XP shows better performance compared to SIDRA-RVS. Hence, we conclude that the XP stellar parameters measured by FS24 provide the better age estimate than the RVS alone. We also explored using FS24 XP stellar parameters and RVS spectra together with both XGBoost and ANN models. However, we could not see any significant increase in the performance. Because the XP spectra are available for many more stars than the RVS spectra, we conclude that the XP spectra alone would provide the valuable age information for more giant stars.
Additionally, we explore the case of training SIDRA-XP model without the [C/Fe] and [N/Fe] information. The model still performs effectively, albeit with slightly greater uncertainty compared to when the [C/Fe] and [N/Fe] parameters were included. The standard deviation of the residuals for the test data around $\mathrm{log}_{10}(\tau_{\mathrm{BINGO}}$ $\mathrm{[Gyr]})=1$ is approximately 0.068 dex. Consequently, there is approximately 6% improvement when incorporating the [C/N] data. We compare the SIDRA-XP ages with the stellar ages from Kordopatis et al. (2023) for the same stars. We find that the ages from Kordopatis et al. (2023) do not display the expected correlation with [Fe/H] or [$\alpha$/M], which is typically seen with age as shown in the next section. Consequently, the ages from Kordopatis et al. (2023) appear to be less reliable than the SIDRA-XP ages. However, we stress that this success is likely due to the carefully selected sample of the giant stars in our study following Ciucă et al. (2021). Our SIDRA-XP is only applicable for the carefully selected populations of the high-mass RC and RGB stars. In the next section, we demonstrate the power of the age estimates of the large number of the giant stars from the Gaia XP spectra.
@@ 5 Chronological map of the Galactic disc with SIDRA-XP
We use the trained model, SIDRA-XP, described in Sec. 4.1 for analysing the remainder of the FS24 dataset. From this dataset, we select stars that meet the following criteria: $T_{\mathrm{eff_{u}nc}}<69.1$ K, log $g_{\mathrm{unc}}<0.14$ dex, $[\mathrm{Fe}/\mathrm{H}]_{\mathrm{unc}}<0.068$ dex, $54004000<T_{\mathrm{eff}}<$ 5400 K and $3.51.5<\log g<$ 3.5. To exclude low-mass RC stars, we further exclude the stars whose $\log(T_{\mathrm{eff}}[\mathrm{K}])>3.675$ and $\log g<$2.6 dex. In cases where duplicate entries exist, the data with the smaller $T_{\mathrm{eff_{u}nc}}$ is selected.
@@
We further apply the astrometry quality cut for the classified stars, selecting the stars with Renormalised Unit Weight Error, $\texttt{RUWE}<1.4$ (Lindegren et al., 2021b) and low parallax error $\mathtt{parallax\_over\_error}>5$. We end up with a total of 2,218,154 stars to use for our SIDRA-XP model.
@@
For our data analysis, the distances to the stars are obtained simply by inverse of the Gaia’s parallax measurements (Lindegren et al., 2021a). We also use the zero-point correction of the parallax as suggested by Lindegren et al. (2021b). We then assume that the Sun’s height with respect to the midplane is $z_{\odot}=0.0208$ kpc (Bennett & Bovy, 2019) and assume $R_{0}=8.275$ kpc (GRAVITY Collaboration et al., 2021).
@@
In this section, we examine the evolution of metallicity, $[\mathrm{Fe/H}]$, and $\alpha$-abundances, $[\mathrm{\alpha/M}]$, for the 2,218,154 stars, which were obtained from FS24 XP stellar data as a function of age, derived using SIDRA-XP, $\tau_{\mathrm{SIDRA}}$. Fig. 8 illustrates $[\mathrm{Fe/H}]$ as a function of $\tau_{\mathrm{SIDRA}}$, with colours representing $[\mathrm{\alpha/M}]$ (top panel), $[\mathrm{\alpha/M}]$ as a function of $\tau_{\mathrm{SIDRA}}$, coloured by metallicity, $[\mathrm{Fe/H}]$, (middle panel) and the distribution of $[\mathrm{\alpha/M}]$ and [Fe/H], coloured by age (bottom panel) across the radial extent of the Galactic disc. In the top and middle panels of Fig. 8, the observed deficiency of stars around $\tau_{\mathrm{SIDRA}}=2.5$ Gyr is due to the selection of high-mass RC stars and a significantly lower number of RGB stars younger than $\sim 3$ Gyr.
@@
The top panel examines the age-metallicity relationship coloured with $[\mathrm{\alpha/M}]$ for SIDRA-XP’s age results. The dataset across the radial extent of the Galactic disc show a flat age-[Fe/H] relation up to approximately 11 Gyr for the younger, low-$[\mathrm{\alpha/M}]$ disc stars, and the expected decrease in [Fe/H] for older, high-$[\mathrm{\alpha/M}]$ stars is also observed in SIDRA-XP’s age results. Our findings qualitatively aligns with the trend seen in the APOGEE data with astroseismic ages (see also Silva Aguirre et al., 2018; Mackereth et al., 2019; Miglio et al., 2021).
As outlined by Ciucă et al. (2024), the Galactic disc evolution can be divided into three phases highlighted in the upper panels of Fig. 8. The first phase, which Ciucă et al. (2024) calls Babi, represents an old, metal-poor disc population older than $\tau_{\mathrm{SIDRA}}>12$ Gyr with a metallicity [M/H]$\simeq-0.4$. The second phase, known as the Great Galactic Starburst (GGS) by Ciucă et al. (2024), occurs between $129<\tau_{\mathrm{SIDRA}}<$ 12 Gyr and is characterised by a rapid increase in metallicity, likely due to a gas-rich merger event of the Gaia-Sausage Enceladus (GSE Belokurov et al., 2018; Helmi et al., 2018). During this phase, older stars show lower [Fe/H] and higher $[\mathrm{\alpha/M}]$ ratios, transitioning to higher [Fe/H] and lower $[\mathrm{\alpha/M}]$ as the thin disc begins to form. The third phase is the thin disc formation, continuing from 9 Gyr ago, where younger stars exhibit higher metallicities and lower $[\mathrm{\alpha/M}]$.
@@
The density plot of Fig. 9 also reveals a noticeable feature around $\tau_{\mathrm{SIDRA}}$ $\sim$ 7 Gyr with [Fe/H]$\simeq-0.4$. Interestingly, this population also has a higher $[\mathrm{\alpha/M}]$ and creating a sequence of low-[Fe/H] from [Fe/H]$\simeq-0.4$ at $\sim 7$ Gyr to [Fe/H]$\simeq-0.1$ at $\sim$ 2 Gyr in the $\tau_{\mathrm{SIDRA}}$-[Fe/H] plot, $[\mathrm{\alpha/M}]$ $\simeq 0.1$ at $\sim$ 7 Gyr to $[\mathrm{\alpha/M}]$ $\simeq 0.05$ at $\sim$ 2 Gyr in the $\tau_{\mathrm{SIDRA}}$-$[\mathrm{\alpha/M}]$ plot in Figs. 8 and 9. The creation of new stars with low [M/H] and high $[\mathrm{\alpha/M}]$ ratios is a sign of a gas-rich merger, as it leads to the dilution of [M/H] and promotes star formation followed by enrichment from Type II supernovae, which increases $[\mathrm{\alpha/M}]$ as demonstrated in Brook et al. (2007). These observations suggest that approximately 7 Gyr ago, a satellite galaxy with a significant amount of gas, had a close encounter with the Milky Way gas disc, which triggered a burst of star formation, and introduced metal-poorer gas that later formed new stars. Interestingly, this low [M/H] population around $\tau_{\mathrm{SIDRA}}$ $\sim$ 7 Gyr is more pronounced in the outer disc, as also indicated in Das et al. (2020). Also, this time coincides with the expected first infall time of the Sagittarius dwarf galaxy (Ruiz-Lara et al., 2020). This could be the indication that the Sagitarrius galaxy was large enough and the first interaction of the Sagitarrius was a gas-rich encounter with the outer edge of the disc.
The middle panel of Fig. 8 illustrates the age-$[\mathrm{\alpha/M}]$ relationship coloured with metallicity for SIDRA-XP’s age results. The high-$[\mathrm{\alpha/M}]$ ‘cluster’ separates clearly from the low-$[\mathrm{\alpha/M}]$ ‘cluster’ in the age-$[\mathrm{\alpha/M}]$ space at $[\mathrm{\alpha/M}]$ $\sim 0.15$ dex. Most high-$[\mathrm{\alpha/M}]$ stars ($[\mathrm{\alpha/M}]$ > $0.15$ dex) are generally older and more metal-poor ($[\mathrm{Fe/H}]<-0.2$ dex) compared to the relatively younger low-$[\mathrm{\alpha/M}]$ population which is shown across all Galactic radial ranges.
The bottom panel of Fig. 8 shows the distribution of [Fe/H] and $[\mathrm{\alpha/M}]$ of our disc stars coloured by age, $\tau_{\mathrm{SIDRA}}$, across different radial ranges of the Galactic disc. In this panel, there is a clear trend where younger stars tend to have higher [Fe/H], while older stars are more metal-poor. The SIDRA-XP ages also display a distinction between high-$[\mathrm{\alpha/M}]$ and low-$[\mathrm{\alpha/M}]$ stars across different ages. Older stars tend to have higher $[\mathrm{\alpha/M}]$ abundances, while younger stars show lower $[\mathrm{\alpha/M}]$ abundances. Overall, SIDRA-XP captures the overall age trend from high-$[\mathrm{\alpha/M}]$ and low-[Fe/H] thick disc populations to low-$[\mathrm{\alpha/M}]$ and high-[Fe/H] thin disc population.
For low-[$\alpha$/Fe] and high-[Fe/H] thin disc population generally shows a broad but flat [Fe/H] distribution in the age-[Fe/H] relation (upper panels in Figs.  8 and 9). Interestingly, in addition to the negative age-[Fe/H] sequence for lower [Fe/H] stars as discussed above, there is a positive age-[Fe/H] sequence for higher [Fe/H] from [Fe/H]$\sim$0.2 at $\tau_{\mathrm{SIDRA}}\sim 9$ Gyr to [Fe/H]$\sim$0.1 at $\tau_{\mathrm{SIDRA}}\sim 2$ Gyr as visible in Fig. 9. This positive trend of the age-[Fe/H] relation is likely a driver of the higher SHAP value (older age) for the higher [Fe/H] feature in Fig. 7 and the lower SHAP value (younger age) for the lower [Fe/H] feature. This trend is also seen in previous literature. For example, Xiang & Rix (2022) interpret this feature as a consequence of radial migration, where fewer young, metal-rich stars born in the inner disc migrate to the outer disc. However, we observe this negative age-[Fe/H] sequence in the wide range of the radius. This may require an alternative scenario to explain.
@@
Figure 8: $[\mathrm{Fe/H}]$ as a function of stellar age, $\tau_{\mathrm{SIDRA}}$, coloured by $[\mathrm{\alpha/M}]$ (top panel), $[\mathrm{\alpha/M}]$ as a function of stellar age, $\tau_{\mathrm{SIDRA}}$, coloured by metallicity, $[\mathrm{Fe/H}]$ (middle panel) and the distribution of $[\mathrm{\alpha/M}]$ and [M/H], coloured by age, $\tau_{\mathrm{SIDRA}}$ (bottom panel) across the radial extent of the Galactic disc.
Figure 9: The density plot of $[\mathrm{Fe/H}]$ as a function of stellar age, $\tau_{\mathrm{SIDRA}}$, (top panel), $[\mathrm{\alpha/M}]$ as a function of stellar age, $\tau_{\mathrm{SIDRA}}$, (middle panel), and the distribution of $[\mathrm{\alpha/M}]$ and [Fe/H] (bottom panel) across the radial extent of the Galactic disc .

### v2

@@ 3.2 Age inference from RVS spectra using SIDRA-RVS
Fig. 4 is a SHapley Additive exPlanations (SHAP, Lundberg and Lee, 2017) bee-swarm plot of our model’s results shown in Fig. 2 and demonstrates the relative importance of the flux values at the different wavelength input features in the RVS spectra and how they impact the stellar-age model output. Fig. 4 shows the distribution of the impact across the top 10 high-impact wavelength pixel values. This visualisation is important for understanding our SIDRA-RVS model’s behaviour and the significance of different elemental lines. The x-axis represents the SHAP values, which indicate the impact of each wavelength input feature (normalised flux) on the prediction. SHAP values to the right of the centerline (positive values) suggest that the corresponding input feature contributes to a higher predicted age, whereas values to the left (negative values) indicate that the input feature results in a lower predicted age. The colour of the dots represents the relative values of the input features, with red indicating higher values and blue indicating lower values. Thus, if blue dots appear on the right side, it suggests that lower input feature values correspond to higher output values. The y-axis represents the features sorted by their importance from top to bottom. Here, the importance of the input feature is measured as the summation of the absolute values of SHAP value for all the data. The top elemental features, TiI: 869.47 nm, TiI: 869.46 nm (2nd), and TiI: 869.48 nm (6th), belong to the same absorption line. Similarly, TiI: 852.06 nm (3rd), TiI: 852.05 nm (8th), and TiI: 852.04 nm (10th) are part of the same line. Additionally, NI: 868.53 nm (4th) and NI: 858.54 nm (5th) are also in the same absorption line, making it the third most important line. Blue dots on the high SHAP value side indicate that a lower flux value at the TiI lines, representing a stronger TiI absorption feature in the spectra, predicts a higher age. On the other hand, red dots on the low SHAP value side suggest that a higher flux value at TiI, representing a weaker TiI absorption feature in the spectra, predicts a lower age.
Figure 5: Example of a high signal-to-noise ratio (SNR) spectrum of an observed star from Gaia DR3 (DR3 Star ID-1222988540219279360) shown in two wavelength regions: 846-858 nm (top panel) and 858-870 nm (bottom panel). The red dots on the spectrum represent the top 10 highest SHAP values from our SIDRA-RVS model results, overlaid with atomic lines from Recio-Blanco et al. (2023). The width of the atomic lines corresponds to its uncertainty. The vertical blue lines at 865.09 in the bottom panel indicate SiI line identified by Contursi et al. (2021). The stellar parameters, $\log_{10}{(T_{\mathrm{eff}})}$, $\log~g$, $\mathrm{\alpha/M}]$ are obtained from APOGEE.
@@ 4.1 SIDRA-XP’s methodology
The relationships between stellar ages and stellar parameters, such as effective temperature, $T_{\mathrm{eff}}$, surface gravity, log $g$, and various abundance ratios, such as [C/Fe], [Fe/H], [N/Fe] and $\mathrm{\alpha/M}$], can be complex and non-linear. ANNs are well-suited for this task because they can autonomously identify and model complex relationships without the need for explicitly programmed rules. ANNs are computational models inspired by the human brain’s neural networks. They consist of layers of interconnected nodes, or "neurons", which process input data to learn patterns and relationships. In supervised learning, the network’s parameters, such as the weights connecting the neurons, are trained and optimised to accurately reproduce known input-output pairs from a training set. Once trained, the ANN can efficiently predict the ages of the other stars based on their observed properties, using significantly less computational effort than the training process.
We split the data 80% training data and 20% testing data. To mitigate the data imbalance between young and old stars, we apply the same random-sampling data-augmentation technique described in Sec. 3.1 to address the underrepresented ages for the training set of the BINGO-FS24. For the data augmentation, we sample $T_{\mathrm{eff}}$, log $g$, [C/Fe], [Fe/H], [N/Fe], $\alpha$/Fe] and age from the original star data using the random Gaussian based on the uncertainties in FS24 and BINGO.
@@ 4.2 Age inference from XP stellar parameters using SIDRA-XP
Figure 6: SIDRA-XP age predictions, $\mathrm{log}_{10}(\tau_{\mathrm{SIDRA}}$ $\mathrm{[Gyr]})$, versus the target BINGO age estimation in $\mathrm{log}_{10}(\tau_{\mathrm{BINGO}}$ $\mathrm{[Gyr]})$ for the XP parameters testing set colour-coded by [Fe/H] (left panel) and $\mathrm{\alpha/M}$] (right panel) obtained from FS24. The upper panels show the predictions versus the target and the black line indicates the identity line. The lower panels represent the residuals between the SIDRA-XP’s $\mathrm{log}_{10}(\tau_{\mathrm{SIDRA}}$ $\mathrm{[Gyr]})$ prediction and BINGO’s true $\mathrm{log}_{10}(\tau_{\mathrm{BINGO}}$ $\mathrm{[Gyr]})$ denoted as $\mathrm{log}_{10}(\frac{\tau_{\mathrm{SIDRA}}}{\tau_{\mathrm{BINGO}}})$. The black filled circles and vertical error bars indicate the mean and the standard deviation of the residuals at the different $\mathrm{log}_{10}(\tau_{\mathrm{BINGO}}$ $\mathrm{[Gyr]})$ bins, respectively.
@@
The colours of the dots in Fig. 6 indicate [Fe/H] (left panel) and $\mathrm{\alpha/M}$] (right panel). The figure shows that there is scatter in [Fe/H] and $\alpha$/M] for the stars with similar age. This means that the SIDRA-XP model does not exclusively capture the correlations between the $\mathrm{\alpha/M}$] or [Fe/H] and stellar age when using XP parameters, ensuring that the model’s learning process goes beyond these specific trends, capturing a broader range of stellar characteristics.
@@
Fig. 7 is a SHAP bee-swarm plot that illustrates the contribution of various input features to the predicted ages of stars using SIDRA-XP. Each dot represents an individual star in the test set result as shown in Fig. 6. The x-axis indicates the SHAP value, representing the impact of each feature on the model’s age prediction for a given star. The features are listed on the y-axis in order from the highest to lowest importance of the input features, which includes effective temperature, $T_{\mathrm{eff}}$, surface gravity, log $g$, and different abundance ratios, such as [C/Fe], [Fe/H], [N/Fe] and $\mathrm{\alpha/M}$]. The distribution and spread of the dots indicate how strongly each XP stellar parameter feature influences the stellar age prediction, with dots further from zero having a more significant impact. The alignment of the dots along the x-axis shows whether the features have a positive or negative contribution to the predicted age. In the case of $T_{\mathrm{eff}}$, the bluer dots in the high SHAP values for $T_{\mathrm{eff}}$ indicate that low $T_{\mathrm{eff}}$ stars push the age prediction higher and the redder dots in the lower SHAP values mean that high $T_{\mathrm{eff}}$ stars push the age prediction to lower. This trend is consistent with the right panel of Fig. 1 and what is found for SIDRA-RVS in Section 3.2.
@@
The [Fe/H] trend of Fig. 7 is puzzling. It shows the higher [Fe/H] leading to older age. However, there are some stars with blue (low [Fe/H]) in the positive SHAP value (older age). Hence, there seems to be a mixed trend of age-[Fe/H] relation. This likely reflects the positive trend of age-[Fe/H] relation in thin disc stars whose age is less than about 8 Gyr, which is also discussed in the next section (see Fig. 9). On the other hand, the $\alpha$/M] shows a clear trend of higher $\alpha$/M] leading to older age. This is consistent with the known age-$\alpha$/M] relation.
Our results demonstrate that the stellar parameters and [C/N] abundances derived from the XP spectra by FS24 provide reliable stellar age information for the giant stars. Unlike SIDRA-RVS, SIDRA-XP learned age-[C/N] relation in addition to the age dependence on $T_{\mathrm{eff}}$, $\log~g$ and $\alpha$/M]. Also, SIDRA-XP shows better performance compared to SIDRA-RVS. Hence, we conclude that the XP stellar parameters measured by FS24 provide the better age estimate than the RVS alone. We also explored using FS24 XP stellar parameters and RVS spectra together with both XGBoost and ANN models. However, we could not see any significant increase in the performance. Because the XP spectra are available for many more stars than the RVS spectra, we conclude that the XP spectra alone would provide the valuable age information for more giant stars.
Additionally, we explore the case of training SIDRA-XP model without the [C/Fe] and [N/Fe] information. The model still performs effectively, albeit with slightly greater uncertainty compared to when the [C/Fe] and [N/Fe] parameters were included. The standard deviation of the residuals for the test data around $\mathrm{log}_{10}(\tau_{\mathrm{BINGO}}$ $\mathrm{[Gyr]})=1$ is approximately 0.068 dex. Consequently, there is approximately 6% improvement when incorporating the [C/N] data. We compare the SIDRA-XP ages with the stellar ages from Kordopatis et al. (2023) for the same stars. We find that the ages from Kordopatis et al. (2023) do not display the expected correlation with [Fe/H] or $\alpha$/M], which is typically seen with age as shown in the next section. Consequently, the ages from Kordopatis et al. (2023) appear to be less reliable than the SIDRA-XP ages. However, we stress that this success is likely due to the carefully selected sample of the giant stars in our study following Ciucă et al. (2021). Our SIDRA-XP is only applicable for the carefully selected populations of the high-mass RC and RGB stars. In the next section, we demonstrate the power of the age estimates of the large number of the giant stars from the Gaia XP spectra.
@@ 5 Chronological map of the Galactic disc with SIDRA-XP
As detailed in Appendix C, we use the trained model, SIDRA-XP, described in Sec. 4.1 for analysing the remainder of the FS24 dataset. As detailed in Appendix C, this yields performance consistent with the model trained on the Miglio et al. (2021) labels. From the FS24 dataset, we select stars that meet the following criteria: $T_{\mathrm{eff_{unc}}}<69.1$ K, log $g_{\mathrm{unc}}<0.14$ dex, $\mathrm{Fe}/\mathrm{H}]_{\mathrm{unc}}<0.068$ dex, $4000<T_{\mathrm{eff}}<5400$ K and $1.5<\log g<3.5$. To exclude low-mass RC stars, we further exclude the stars whose $\log(T_{\mathrm{eff}}[\mathrm{K}])>3.675$ and $\log g<$2.6 dex. In cases where duplicate entries exist, the data with the smaller $T_{\mathrm{eff_{unc}}}$ is selected.
@@
We further apply the astrometry quality cut for the classified stars, selecting the stars with Renormalised Unit Weight Error, $\texttt{RUWE}<1.4$ (Lindegren et al., 2021a) and low parallax error $\mathtt{parallax\_over\_error}>5$. We end up with a total of 2,218,154 stars to use for our SIDRA-XP model.
@@
For our data analysis, the distances to the stars are obtained simply by inverse of the Gaia’s parallax measurements (Lindegren et al., 2021b). We also use the zero-point correction of the parallax as suggested by Lindegren et al. (2021a). We then assume that the Sun’s height with respect to the midplane is $z_{\odot}=0.0208$ kpc (Bennett and Bovy, 2019) and assume $R_{0}=8.275$ kpc (GRAVITY Collaboration et al., 2021).
@@
In this section, we examine the evolution of metallicity, $\mathrm{Fe/H}]$, and $\alpha$-abundances, $\mathrm{\alpha/M}]$, for the 2,218,154 stars, which were obtained from FS24 XP stellar data as a function of age, derived using SIDRA-XP, $\tau_{\mathrm{SIDRA}}$. Fig. 8 illustrates $\mathrm{Fe/H}]$ as a function of $\tau_{\mathrm{SIDRA}}$, with colours representing $\mathrm{\alpha/M}]$ (top panel), $\mathrm{\alpha/M}]$ as a function of $\tau_{\mathrm{SIDRA}}$, coloured by metallicity, $\mathrm{Fe/H}]$, (middle panel) and the distribution of $\mathrm{\alpha/M}]$ and [Fe/H], coloured by age (bottom panel) across the radial extent of the Galactic disc. In the top and middle panels of Fig. 8, the observed deficiency of stars around $\tau_{\mathrm{SIDRA}}=2.5$ Gyr is due to the selection of high-mass RC stars and a significantly lower number of RGB stars younger than $\sim 3$ Gyr.
@@
The top panel examines the age-metallicity relationship coloured with $\mathrm{\alpha/M}]$ for SIDRA-XP’s age results. The dataset across the radial extent of the Galactic disc show a flat age-[Fe/H] relation up to approximately 11 Gyr for the younger, low-$\mathrm{\alpha/M}]$ disc stars, and the expected decrease in [Fe/H] for older, high-$\mathrm{\alpha/M}]$ stars is also observed in SIDRA-XP’s age results. Our findings qualitatively aligns with the trend seen in the APOGEE data with astroseismic ages (see also Silva Aguirre et al., 2018; Mackereth et al., 2019; Miglio et al., 2021).
[…]
@@
[…]
The middle panel of Fig. 8 illustrates the age-$\mathrm{\alpha/M}]$ relationship coloured with metallicity for SIDRA-XP’s age results. The high-$\mathrm{\alpha/M}]$ ‘cluster’ separates clearly from the low-$\mathrm{\alpha/M}]$ ‘cluster’ in the age-$\mathrm{\alpha/M}]$ space at $\mathrm{\alpha/M}]$ $\sim 0.15$ dex. Most high-$\mathrm{\alpha/M}]$ stars ($\mathrm{\alpha/M}]$ > $0.15$ dex) are generally older and more metal-poor ($\mathrm{Fe/H}]<-0.2$ dex) compared to the relatively younger low-$\mathrm{\alpha/M}]$ population which is shown across all Galactic radial ranges.
The bottom panel of Fig. 8 shows the distribution of [Fe/H] and $\mathrm{\alpha/M}]$ of our disc stars coloured by age, $\tau_{\mathrm{SIDRA}}$, across different radial ranges of the Galactic disc. In this panel, there is a clear trend where younger stars tend to have higher [Fe/H], while older stars are more metal-poor. The SIDRA-XP ages also display a distinction between high-$\mathrm{\alpha/M}]$ and low-$\mathrm{\alpha/M}]$ stars across different ages. Older stars tend to have higher $\mathrm{\alpha/M}]$ abundances, while younger stars show lower $\mathrm{\alpha/M}]$ abundances. Overall, SIDRA-XP captures the overall age trend from high-$\mathrm{\alpha/M}]$ and low-[Fe/H] thick disc populations to low-$\mathrm{\alpha/M}]$ and high-[Fe/H] thin disc population.
[…]
@@
Figure 8: $\mathrm{Fe/H}]$ as a function of stellar age, $\tau_{\mathrm{SIDRA}}$, coloured by $\mathrm{\alpha/M}]$ (top panel), $\mathrm{\alpha/M}]$ as a function of stellar age, $\tau_{\mathrm{SIDRA}}$, coloured by metallicity, $\mathrm{Fe/H}]$ (middle panel) and the distribution of $\mathrm{\alpha/M}]$ and $\mathrm{Fe/H}]$, coloured by age, $\tau_{\mathrm{SIDRA}}$ (bottom panel) across the radial extent of the Galactic disc.
Figure 9: The density plot of $\mathrm{Fe/H}]$ as a function of stellar age, $\tau_{\mathrm{SIDRA}}$, (top panel), $\mathrm{\alpha/M}]$ as a function of stellar age, $\tau_{\mathrm{SIDRA}}$, (middle panel), and the distribution of $\mathrm{\alpha/M}]$ and [Fe/H] (bottom panel) across the radial extent of the Galactic disc .